A pail of water of mass 5.0 kg is suspended from a weighing scale hanging from the ceiling of a lift.What is the reading of the scale when the lift moves downward at a deceleration of 4ms-2

1
Expert's answer
2021-02-25T12:36:09-0500
"N-mg=-ma,""N=m(g-a)=5.0\ kg\cdot(9.8\ \dfrac{m}{s^2}-4.0\ \dfrac{m}{s^2})=29\ N."